OpenStudy (anonymous):

Trigonometry Help Prove the Identity (sinx - cosx)/(secx - cscx) = (sin2x)/2 I changed to cos-sin stuff and got: (sinx - cosx)/((1/cosx)-(1/sinx)) but here i get stuck once again, idk where to go from here..

ganeshie8 (ganeshie8):

multiply the numerator and denominator by \(\sin x \cos x\)

OpenStudy (anonymous):



ganeshie8 (ganeshie8):

multiply the numerator and denominator by \(\sin x \cos x\)

OpenStudy (anonymous):

but gane, where is the sinx cosx even from? i just see sinx-cosx

OpenStudy (anonymous):

so the sinxcosx is from the right side of the equation, i am just using it on the left?

ganeshie8 (ganeshie8):

nope, lets work it again from scratch

ganeshie8 (ganeshie8):

we take left hand side, massage it such that we get the right hand side

OpenStudy (raden):

secx - cscx = 1/c - 1/s = (s - c)/sc

ganeshie8 (ganeshie8):

so we start with left hand side : \[\large \dfrac{\sin x - \cos x }{\sec x - \csc x} \]

ganeshie8 (ganeshie8):

your earlier first step was correct, change sec and csc to cos and sin in the denominator

OpenStudy (anonymous):

alright and we change everything to sin/cos

ganeshie8 (ganeshie8):

yes

ganeshie8 (ganeshie8):

\[\large \dfrac{\sin x - \cos x }{\sec x - \csc x} = \dfrac{\sin x - \cos x }{\frac{1}{\cos x}- \frac{1}{\sin x}} \]

ganeshie8 (ganeshie8):

Now stare at the denominator fractions

ganeshie8 (ganeshie8):

can you combine them by getting a common denominator ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

whats the LCM of \(\cos x \), \(\sin x\) ?

OpenStudy (anonymous):

well we want to get it to sinxcosx correct?

ganeshie8 (ganeshie8):

yes

OpenStudy (anonymous):

so it would be..\[\frac{ sinx-cox }{ sinxcosx }\]

ganeshie8 (ganeshie8):

Correct! thats the denominator

OpenStudy (anonymous):

alright alright, next we want to get rid of the denominator's denominator?

OpenStudy (anonymous):

well more like we dont want the denominator to be a fraction

ganeshie8 (ganeshie8):

\[\large \dfrac{\sin x - \cos x }{\sec x - \csc x} = \dfrac{\sin x - \cos x }{\frac{1}{\cos x}- \frac{1}{\sin x}}\] \[ \large = \dfrac{\sin x - \cos x}{\frac{\sin x - \cos x}{\sin x \cos x}} \]

ganeshie8 (ganeshie8):

yes we don't want fractions

ganeshie8 (ganeshie8):

can we cancel sinx-cosx ?

OpenStudy (anonymous):

alright, im not really too sure how to do that, im not good with fractions.

ganeshie8 (ganeshie8):

\[\large \dfrac{\sin x - \cos x }{\sec x - \csc x} = \dfrac{\sin x - \cos x }{\frac{1}{\cos x}- \frac{1}{\sin x}}\] \[ \large = \dfrac{\sin x - \cos x}{\frac{\sin x - \cos x}{\sin x \cos x}} \] \[ \large = \dfrac{1}{\frac{1}{\sin x \cos x}} \]

OpenStudy (anonymous):

oh um..yes? since its in the numerator too?

ganeshie8 (ganeshie8):

you got it !

ganeshie8 (ganeshie8):

now multiply numerator and denominator to get rid of the fraction in the bottom

ganeshie8 (ganeshie8):

\[ \large = \dfrac{1}{\frac{1}{\sin x \cos x}} \] \[ \large = \dfrac{\color{red}{\sin x \cos x} }{\color{red}{\sin x \cos x} \frac{1}{\sin x \cos x}} \]

ganeshie8 (ganeshie8):

in the denominator sinxcosx cancels out, leaving you with just sinxcosx in the numerator

ganeshie8 (ganeshie8):

\[ \large = \dfrac{1}{\frac{1}{\sin x \cos x}} \] \[ \large = \dfrac{\color{red}{\sin x \cos x} }{\color{red}{\sin x \cos x} \frac{1}{\sin x \cos x}} \] \[\large = \color{red}{\sin x \cos x}\]

OpenStudy (anonymous):

ooh ok since it would just be sinxcosx/1 it makes it just sinxcosx i see now

OpenStudy (anonymous):

but how does that become sin2x/2..that makes no sense to me

ganeshie8 (ganeshie8):

yes next you need to recall the identity : \(\sin 2x = 2 \sin x \cos x\)

OpenStudy (anonymous):

ok i see that, under the Double Angle Formulas

ganeshie8 (ganeshie8):

yes thats the name ! using that identity, conclude that \(\large \color{red}{\sin x \cos x = \dfrac{\sin 2x}{2}}\)

OpenStudy (anonymous):

im sorry but i just dont understand the last part, with the sin 2x = 2sinx cosx. i just dont understand how that makes sinxcosx into sin2x/2

ganeshie8 (ganeshie8):

we have this identity : \(\large \sin 2x = 2\sin x \cos x\)

ganeshie8 (ganeshie8):

divide 2 both sides, what do you get ?

OpenStudy (anonymous):

oh lol i see now

ganeshie8 (ganeshie8):

good :) \(\large \sin x \cos x = \dfrac{2}{2} \sin x \cos x = \dfrac{1}{2} \sin (2x)\)

OpenStudy (anonymous):

ugh everyone on here is better than my professor at explaining everything lol. Thank you so much ganeshie8

OpenStudy (ikram002p):

huh @nsellers23 u wont find better than ganesh in explaining :P

OpenStudy (anonymous):

haha i dont doubt that one bit ikram

ganeshie8 (ganeshie8):

lol i am flattered ! ty for the good words :) good luck !!

OpenStudy (anonymous):

just 5 more problems to go, ill see you guys in a bit if i cant get them